OpenStudy (anonymous):

write slope-intercept form (-4,4)(4,-2)

OpenStudy (anonymous):

I'm inclined to write in: y=-4x+4 and y=4x+(-2), but do I have to solve also?

OpenStudy (anonymous):

You must first find your slope. To do that remember \[(y _{2} -y_{1})/(x_{2}-x_{1})\] This gives you (-2-4)/(4-(-4)) = -6/8 = -3/4

OpenStudy (anonymous):

so, as far as the form y-mx+b.....I can plug in y=-3/4x+?

OpenStudy (anonymous):

After that use point slope for to get y - 4 = -3/4(x-(-4)) - > y -4 = -3/4x - 3 -> y = -3/4x +1

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

You need to either use point-slope form or determine your y intercept after finding the slope before you can write it in slope-intercept form.. The above is the answer.

OpenStudy (anonymous):

thank you decemberchild